# Padé Approximant Neural Networks for Enhanced Electric Motor Fault Diagnosis Using Vibration and Acoustic Data

### v1

@@ 3.2 Data Preparation
All channels of vibration and audio data were segmented into fixed-length windows of 1000 time-domain samples without overlap to ensure uniformity across the dataset. Thus, a total of 420 samples were extracted per channel from a single recording. The segmentation process divides the continuous time series data into smaller, non-overlapping segments of 1000 samples to balance temporal resolution and model complexity. Each segment is then normalized separately. Normalization is performed using the min-max normalization technique, where each data point $x_{i}$ in a given segment is scaled based on the minimum and maximum values of that segment. The normalization equation can be written as:
@@
Figure 4: Examples of segmented and normalized waveforms for 45 Hz input frequency: (a) accelerometer 1, (b) microphone.
@@ 3.3 Model Architecture and Training Strategy
Table 3: Detailed architecture of the proposed 1D PadéNet model is given. Activations in the 1D Padé layers (tanh and LeakyReLU with negative slope coefficient 0.01) were evaluated separately and not explicitly shown.
@@ 15 Flatten (None, 224)
17 Dense (None, 64) Activation = tanh
@@ 18 Dense (None, 8) Activation = Softmax
Following each 1D Padé layer, the activation function applied is either the tanh or LeakyReLU with a negative slope of 0.01. We evaluate the performance of both activation functions separately for 1D CNN and PadéNet. However, in 1D Self-ONNs, we are restricted to using tanh to prevent excessively large activations due to the increasing powers of the input, thus mitigating the risk of the exploding gradients. For all models, a 1D max-pooling layer with a pool size of 2 was applied after each block for spatial downsampling, reducing the dimensionality of the feature maps while preserving the most crucial information.
Each of the seven 1D Padé layers was configured with 32 filters, a kernel size of 7, padding=’same’, and incorporated an L2 kernel regularizer ($\lambda=10^{-4}$) for weight decay. Following these layers, a Flatten layer was used to reshape the feature maps into a vector. It was followed by a dense layer with 64 neurons and a tanh activation, which processes the flattened features. The model ends with a Softmax output layer with 8 neurons for multi-class classification, assigning probabilities to each of the 8 classes. To prevent overfitting, a Dropout layer with a rate of 0.25 was applied before the dense layers during training.
@@ 3.4 Evaluation Metrics
[…]
@@ 4 Results and Discussion
In this section, we present a comprehensive evaluation of the proposed 1D PadéNet compared to 1D CNN and Self-ONN using vibration and acoustic data. Performance metrics were computed for each sensor input on 5 independent runs, and the results were averaged, with standard deviations reported to quantify variability, as detailed in Tables 4, 5, 6, and 7. Aggregated confusion matrices over 5 runs (Figures 5, 6, 7, 8) provide further insight into classification performance across fault types.
@@
$P=1,Q=0$ (tanh) 99.03 100.00 99.69 $\pm$ 0.35 99.70 $\pm$ 0.33 99.69 $\pm$ 0.35 99.69 $\pm$ 0.35
$P=1,Q=0$ (LeakyReLU) 99.33 100.00 99.74 $\pm$ 0.22 99.74 $\pm$ 0.22 99.74 $\pm$ 0.22 99.74 $\pm$ 0.22
$P=1,Q=1$ (tanh) 99.85 100.00 99.95 $\pm$ 0.06 99.95 $\pm$ 0.06 99.95 $\pm$ 0.06 99.95 $\pm$ 0.06
$P=1,Q=1$ (LeakyReLU) 99.89 99.96 99.92 $\pm$ 0.03 99.92 $\pm$ 0.03 99.92 $\pm$ 0.03 99.92 $\pm$ 0.03
$P=1,Q=2$ (tanh) 99.85 100.00 99.96 $\pm$ 0.05 99.96 $\pm$ 0.05 99.96 $\pm$ 0.05 99.96 $\pm$ 0.05
$P=1,Q=2$ (LeakyReLU) 99.85 99.96 99.93 $\pm$ 0.04 99.93 $\pm$ 0.04 99.93 $\pm$ 0.04 99.93 $\pm$ 0.04
$P=2,Q=0$ (tanh) 99.81 100.00 99.95 $\pm$ 0.07 99.95 $\pm$ 0.07 99.95 $\pm$ 0.07 99.95 $\pm$ 0.07
$P=2,Q=1$ (tanh) 99.89 100.00 99.96 $\pm$ 0.04 99.96 $\pm$ 0.04 99.96 $\pm$ 0.04 99.96 $\pm$ 0.04
$P=2,Q=1$ (LeakyReLU) 99.52 99.89 99.70 $\pm$ 0.14 99.70 $\pm$ 0.14 99.70 $\pm$ 0.14 99.70 $\pm$ 0.14
$P=3,Q=0$ (tanh) 99.52 99.96 99.84 $\pm$ 0.17 99.85 $\pm$ 0.16 99.84 $\pm$ 0.17 99.84 $\pm$ 0.17
For the first accelerometer, located at the motor’s drive end, the 1D CNN $(P=1,Q=0)$ configuration with tanh activation function achieved an average test accuracy of 99.69% $\pm$ 0.35%. Replacing the activation function with LeakyReLU (with a negative slope of 0.01) across all convolutional layers led to a noticeable improvement, increasing the average accuracy to 99.74% $\pm$ 0.22%. Among the evaluated Self-ONN configurations, the model with $P=2$ and $Q=0$ attained the highest average test accuracy of 99.95% $\pm$ 0.07%, outperforming all CNN-based counterparts. However, the best overall fault diagnosis performance was achieved by the PadéNet model with $P=2$, $Q=1$ and tanh activation, which reached an average test accuracy of 99.96% $\pm$ 0.04% and an F1-score of 99.96% $\pm$ 0.04%. A detailed summary of the test metrics for all 1D CNN, Self-ONN, and PadéNet models evaluated using the first accelerometer data is provided in Table 4.
@@
Figure 5: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=2,Q=0$), and PadéNet ($P=2,Q=1$) models with the first accelerometer data as input. The 1D Self-ONN and PadéNet models use the tanh activation function, while the 1D CNN utilizes LeakyReLU with negative slope of 0.01.
@@
$P=1,Q=0$ (tanh) 92.75 94.57 93.94 $\pm$ 0.64 93.96 $\pm$ 0.63 93.94 $\pm$ 0.64 93.93 $\pm$ 0.64
$P=1,Q=0$ (LeakyReLU) 96.06 97.62 96.81 $\pm$ 0.51 96.83 $\pm$ 0.49 96.81 $\pm$ 0.51 96.80 $\pm$ 0.50
$P=1,Q=1$ (tanh) 93.79 95.35 94.79 $\pm$ 0.53 94.82 $\pm$ 0.54 94.79 $\pm$ 0.53 94.79 $\pm$ 0.53
$P=1,Q=1$ (LeakyReLU) 96.61 97.88 97.25 $\pm$ 0.43 97.26 $\pm$ 0.43 97.25 $\pm$ 0.43 97.25 $\pm$ 0.43
$P=1,Q=2$ (tanh) 95.05 96.28 95.47 $\pm$ 0.43 95.47 $\pm$ 0.43 95.47 $\pm$ 0.43 95.46 $\pm$ 0.43
$P=1,Q=2$ (LeakyReLU) 96.95 97.81 97.54 $\pm$ 0.32 97.55 $\pm$ 0.33 97.54 $\pm$ 0.32 97.54 $\pm$ 0.32
$P=2,Q=0$ (tanh) 95.31 96.24 95.96 $\pm$ 0.33 95.97 $\pm$ 0.32 95.96 $\pm$ 0.33 95.96 $\pm$ 0.33
$P=2,Q=1$ (tanh) 96.35 97.25 96.85 $\pm$ 0.29 96.86 $\pm$ 0.28 96.85 $\pm$ 0.29 96.84 $\pm$ 0.29
$P=2,Q=1$ (LeakyReLU) 97.88 98.59 98.26 $\pm$ 0.26 98.27 $\pm$ 0.26 98.26 $\pm$ 0.26 98.26 $\pm$ 0.26
$P=3,Q=0$ (tanh) 96.13 97.21 96.67 $\pm$ 0.34 96.68 $\pm$ 0.35 96.67 $\pm$ 0.34 96.67 $\pm$ 0.35
Located on the bearing housing near the drive end, the second accelerometer presented a complex diagnostic environment due to potential signal disturbances from bearing dynamics. When the second accelerometer data was used as input, the 1D CNN model ($P=1$, $Q=0$) with tanh activation reached an average test accuracy of 93.94% $\pm$ 0.64% as given in Table 5. Switching to LeakyReLU activation across its convolutional layers significantly enhanced performance, boosting the average accuracy to 96.81% $\pm$ 0.51%, with reduced variability indicating improved consistency across each run. Among Self-ONN models tested, the configuration ($P=3$, $Q=0$) delivered the highest mean accuracy of 96.67% $\pm$ 0.34%. Although the average classification accuracy of the Self-ONN model ($P=3$, $Q=0$) surpassed that of the 1D CNN with tanh activation, switching to LeakyReLU enabled the 1D CNN to achieve higher fault diagnosis accuracy. Thus, it can be concluded that Self-ONNs generally outperform CNNs when tanh is employed as the activation function in both models. However, when LeakyReLU is used in all convolutional layers of a 1D CNN, it can achieve higher fault diagnosis accuracy compared to Self-ONNs utilizing the tanh activation function, potentially due to the vanishing gradient issue commonly associated with tanh. In contrast to other 1D CNN and Self-ONN models, the most effective fault diagnosis for the second accelerometer was achieved by the 1D PadéNet model with $P=2,Q=1$ and LeakyReLU activation, having a mean test accuracy of 98.26% $\pm$ 0.26%, along with an F1-score of 98.26% $\pm$ 0.26%. A complete summary of performance metrics for all tested 1D CNN, Self-ONN, and PadéNet models using the second accelerometer data is available in Table 5.
@@
Figure 6: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=2,Q=1$) models with the second accelerometer data as input. The 1D Self-ONN model uses the tanh activation function, while the 1D CNN and PadéNet utilize LeakyReLU with negative slope of 0.01.
@@
Figure 6 shows the aggregated confusion matrices across 5 independent runs for the 1D CNN $(P=1,Q=0)$, Self-ONN $(P=3,Q=0)$, and PadéNet $(P=2,Q=1)$ models, using data from the second accelerometer as input. For these confusion matrices, the 1D Self-ONN model uses the tanh activation function, while the 1D CNN and PadéNet utilize LeakyReLU with negative slope of 0.01. The 1D CNN $(P=1,Q=0)$ and Self-ONN $(P=3,Q=0)$ models achieve classification accuracies of 94.82% and 93.75%, respectively, in the faulty bearings (FB) class. In comparison, the 1D PadéNet $(P=2,Q=1)$ model delivers a significantly higher accuracy of 96.78% in the same class. Similarly, for the rotor misalignment (RM) class, the 1D CNN and Self-ONN models yield accuracies of 93.63% and 94.70%, respectively, whereas the 1D PadéNet $(P=2,Q=1)$ model once again outperforms them with a superior accuracy of 96.67%. Faulty bearings and rotor misalignment are two of the most critical fault types in electric motors, as they can severely compromise mechanical integrity and lead to costly operational downtimes if not detected early. Accurate classification of these faults is therefore essential for timely maintenance and fault prevention. The enhanced fault diagnosis performance of the 1D PadéNet model is clearly evident across all fault classes in these confusion matrices and reflects its superior capability to distinguish even the most critical fault types.
@@
$P=1,Q=0$ (tanh) 93.12 94.38 93.74 $\pm$ 0.45 93.76 $\pm$ 0.44 93.74 $\pm$ 0.45 93.73 $\pm$ 0.44
$P=1,Q=0$ (LeakyReLU) 96.35 97.40 96.73 $\pm$ 0.40 96.74 $\pm$ 0.40 96.73 $\pm$ 0.40 96.72 $\pm$ 0.40
$P=1,Q=1$ (tanh) 93.34 94.61 94.11 $\pm$ 0.50 94.12 $\pm$ 0.50 94.11 $\pm$ 0.50 94.10 $\pm$ 0.50
$P=1,Q=1$ (LeakyReLU) 96.99 97.62 97.30 $\pm$ 0.28 97.32 $\pm$ 0.28 97.30 $\pm$ 0.28 97.30 $\pm$ 0.28
$P=1,Q=2$ (tanh) 93.68 95.76 94.81 $\pm$ 0.73 94.83 $\pm$ 0.73 94.81 $\pm$ 0.73 94.80 $\pm$ 0.74
$P=1,Q=2$ (LeakyReLU) 97.10 97.88 97.61 $\pm$ 0.27 97.62 $\pm$ 0.27 97.61 $\pm$ 0.27 97.61 $\pm$ 0.27
$P=2,Q=0$ (tanh) 92.63 95.46 94.41 $\pm$ 1.01 94.43 $\pm$ 0.99 94.41 $\pm$ 1.01 94.41 $\pm$ 1.01
$P=2,Q=1$ (tanh) 95.39 96.39 96.00 $\pm$ 0.36 96.03 $\pm$ 0.35 96.00 $\pm$ 0.36 96.00 $\pm$ 0.36
$P=2,Q=1$ (LeakyReLU) 97.28 98.03 97.59 $\pm$ 0.27 97.61 $\pm$ 0.27 97.59 $\pm$ 0.27 97.59 $\pm$ 0.27
$P=3,Q=0$ (tanh) 94.72 96.39 95.83 $\pm$ 0.59 95.87 $\pm$ 0.55 95.83 $\pm$ 0.59 95.83 $\pm$ 0.58
When data from the third accelerometer, positioned on the bearing housing farthest from the drive end, is used as input, the results in Table 6 show that the 1D PadéNet model, particularly with $P=1$, $Q=2$ and LeakyReLU activation, consistently outperforms both the 1D CNN and Self-ONN models across all evaluated metrics. In particular, the 1D PadéNet model achieves the highest average test accuracy of 97.61% $\pm$ 0.27%, surpassing the best-performing configurations of the 1D CNN and Self-ONN models.
@@
Figure 7: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=1,Q=2$) models with the third accelerometer data as input. The 1D Self-ONN model uses the tanh activation function, while the 1D CNN and PadéNet utilize LeakyReLU with negative slope of 0.01.
When acoustic signals captured by the microphone are used as input, 1D PadéNet models continue to demonstrate superior diagnostic capabilities, as shown in Table 7. The best performance was achieved by the configuration with $P=1,Q=2$ and LeakyReLU activation, which yielded the highest average accuracy of 98.33% $\pm$ 0.44%. This result is superior to those obtained using data from the second and third accelerometers, yet slightly underperforms compared to the first accelerometer, which is mounted directly above the drive-end.
@@
$P=1,Q=0$ (tanh) 95.05 96.24 95.56 $\pm$ 0.45 95.58 $\pm$ 0.46 95.56 $\pm$ 0.45 95.55 $\pm$ 0.45
$P=1,Q=0$ (LeakyReLU) 96.80 98.03 97.55 $\pm$ 0.46 97.57 $\pm$ 0.45 97.55 $\pm$ 0.46 97.55 $\pm$ 0.46
$P=1,Q=1$ (tanh) 96.84 97.47 97.13 $\pm$ 0.24 97.14 $\pm$ 0.25 97.13 $\pm$ 0.24 97.13 $\pm$ 0.24
$P=1,Q=1$ (LeakyReLU) 97.58 98.36 97.89 $\pm$ 0.30 97.92 $\pm$ 0.29 97.89 $\pm$ 0.30 97.89 $\pm$ 0.31
$P=1,Q=2$ (tanh) 97.47 97.66 97.58 $\pm$ 0.07 97.59 $\pm$ 0.07 97.58 $\pm$ 0.07 97.58 $\pm$ 0.07
$P=1,Q=2$ (LeakyReLU) 97.81 98.92 98.33 $\pm$ 0.44 98.34 $\pm$ 0.43 98.33 $\pm$ 0.44 98.33 $\pm$ 0.44
$P=2,Q=0$ (tanh) 97.62 98.10 97.89 $\pm$ 0.16 97.89 $\pm$ 0.16 97.89 $\pm$ 0.16 97.89 $\pm$ 0.16
$P=2,Q=1$ (tanh) 97.02 98.62 98.07 $\pm$ 0.55 98.08 $\pm$ 0.54 98.07 $\pm$ 0.55 98.07 $\pm$ 0.55
$P=2,Q=1$ (LeakyReLU) 97.92 98.70 98.30 $\pm$ 0.32 98.31 $\pm$ 0.31 98.30 $\pm$ 0.32 98.30 $\pm$ 0.32
$P=3,Q=0$ (tanh) 97.51 98.40 97.93 $\pm$ 0.32 97.94 $\pm$ 0.32 97.93 $\pm$ 0.32 97.93 $\pm$ 0.32
Figure 8: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=1,Q=2$) models with the acoustic data as input. The 1D Self-ONN model uses the tanh activation function, while the 1D CNN and PadéNet utilize LeakyReLU with negative slope of 0.01.
@@
Table 8: Number of trainable parameters and average inference time for each model.
Model # Params Inf. Time (ms)
$P=1$, $Q=0$ (CNN) 58,376 1.368 ± 0.499
$P=1$, $Q=1$ (PadéNet) 101,608 2.197 ± 0.501
$P=1$, $Q=2$ (PadéNet) 144,840 3.127 ± 0.724
$P=2$, $Q=0$ (Self-ONN) 101,832 2.353 ± 0.622
$P=2$, $Q=1$ (PadéNet) 145,064 3.096 ± 0.530
$P=3$, $Q=0$ (Self-ONN) 145,288 3.195 ± 0.796
The computational demands of the 1D CNN, 1D Self-ONN, and 1D PadéNet models were evaluated through their number of trainable parameters and average inference times, as summarized in Table 8. These metrics are critical for assessing the feasibility of deploying fault diagnosis models in resource-constrained devices, where computational efficiency must be balanced against diagnostic accuracy. The average inference time, along with its standard deviation, was computed by performing 10 forward passes of a single data sample through each model using the TensorFlow framework on an NVIDIA RTX 3070 GPU. The 1D CNN ($P=1$, $Q=0$) model requires the fewest parameters (58,376) and exhibits the fastest inference time (1.368 ms $\pm$ 0.499 ms). However, as previously discussed, it typically delivers lower diagnostic accuracy compared to 1D PadéNet models. Additionally, while increasing the degree of the numerator ($P$) and denominator ($Q$) generally enhances fault diagnosis accuracy, it also results in increased computational complexity for the 1D PadéNet models. Nevertheless, with moderate values of $P$ and $Q$, the increase in trainable parameters remains manageable, and the inference time stays within a few milliseconds, making 1D PadéNets a competitive option for fault diagnosis systems that can be deployed on the edge devices.
The fault diagnosis performance of 1D PadéNets is also compared with other DL-based methods trained on the University of Ottawa’s constant-speed vibration and audio datasets. In [10], a 15-layer 1D CNN-LSTM model (comprising 6 Conv1D and 2 LSTM layers) was proposed for classifying electrical and mechanical faults. The model takes an input window size of 1000 samples, as used in this study, with the same train-validation-test split ratios and all accelerometer inputs. To evaluate the impact of the LSTM layers, two other variations of the model were tested. First, the two LSTM layers were removed, resulting in a 13-layer CNN, referred to as “13-Layers CNN” in Table 9. In the second experiment, the LSTM layers in the 11th and 12th layers were replaced with Conv1D layers, producing a 15-layer CNN. All results are presented in Table 9 for each accelerometer sensor. For accelerometer 1, the 1D PadéNet ($P=2$, $Q=1$) achieves the same average test accuracy as the 1D CNN-LSTM model but with significantly fewer trainable parameters, and outperforms all other CNN models. Moreover, for accelerometers 2 and 3, 1D PadéNet delivers comparable classification accuracy to the CNN-LSTM model, while the 13-layer and 15-layer CNN models fail to match those accuracies. However, it is worth noting that, due to the sequential nature of LSTM layers, the CNN-LSTM model tends to be slower during training and inference, whereas the 1D PadéNet achieves similar performance with greater computational efficiency.

### v2

@@ 3.2 Data Preparation
All channels of vibration and audio data were segmented into fixed-length windows of 1000 time-domain samples without overlap. Thus, a total of 420 samples were extracted per channel from a single recording. The segmentation process divides the continuous time series data into smaller, non-overlapping segments of 1000 samples to balance temporal resolution and model complexity. Each segment is then normalized separately. Normalization is performed using the min-max normalization method, where each data point $x_{i}$ in a given segment is scaled based on the minimum and maximum values of that segment. The normalization equation can be written as:
@@
Figure 4: Examples of segmented and normalized waveforms for 45 Hz input frequency: (a) Accelerometer-1, (b) microphone.
@@ 3.3 Model Architecture and Training Strategy
Table 3: Detailed architecture of the proposed 1D PadéNet model is given. Activations in the 1D Padé layers ($\tanh$ and $\mathrm{LeakyReLU}$ with negative slope coefficient 0.01) were evaluated separately and not explicitly shown.
@@ 15 Flatten (None, 224)
17 Dense (None, 64) Activation = $\tanh$
@@ 18 Dense (None, 8) Activation = Softmax
Following each 1D Padé layer, the activation function applied is either the $\tanh$ or $\mathrm{LeakyReLU}$ with a negative slope of 0.01. We evaluate the performance of both activation functions separately for 1D CNN and PadéNet. However, in 1D Self-ONNs, we are restricted to using $\tanh$ to prevent excessively large activations due to the increasing powers of the input, thus mitigating the risk of the exploding gradients. For all models, a 1D max-pooling layer with a pool size of 2 was applied after each block for spatial downsampling, reducing the dimensionality of the feature maps while preserving the most crucial information.
Each of the seven 1D Padé layers was configured with 32 filters, a kernel size of 7, padding=’same’, and incorporated an L2 kernel regularizer ($\lambda=10^{-4}$) for weight decay. Following these layers, a Flatten layer was used to reshape the feature maps into a vector. It was followed by a dense layer with 64 neurons and a $\tanh$ activation, which processes the flattened features. The model ends with a Softmax output layer with 8 neurons for multi-class classification, assigning probabilities to each of the 8 classes. To prevent overfitting, a Dropout layer with a rate of 0.25 was applied before the dense layers during training.
@@ 3.4 Evaluation Metrics
$\text{F1\text{-}Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (22)
@@ 4 Results and Discussion
This section presents the fault diagnosis performance of Padé-based models, along with CNNs and Self-ONNs, on the constant-speed vibration and acoustic datasets from Ottawa University [37]. The analysis further evaluates robustness under varying noise levels, examines the sensitivity of Padé layers to different Padé orders and activation functions, and compares the corresponding computational complexities.
4.1 Performance on the Ottawa University Vibration and Acoustic Datasets
We first present a comprehensive evaluation of the proposed 1D PadéNet compared to 1D CNN and Self-ONN using noise-free vibration and acoustic data [37]. Performance metrics were computed for each sensor input on 5 independent runs, and the results were averaged, with standard deviations reported to quantify variability, as detailed in Tables 4, 5, 6, and 7. Aggregated confusion matrices over 5 runs (Figures 5, 6, 7, 8) provide further insight into classification performance across fault types.
@@ 4.1 Performance on the Ottawa University Vibration and Acoustic Datasets
[…]
For the first accelerometer, located at the motor’s drive end, the 1D CNN $(P=1,Q=0)$ configuration with $\tanh$ activation function achieved an average test accuracy of 99.69% $\pm$ 0.35%. Replacing the activation function with $\mathrm{LeakyReLU}$ (with a negative slope of 0.01) across all convolutional layers led to a noticeable improvement, increasing the average accuracy to 99.74% $\pm$ 0.22%. Among the evaluated Self-ONN configurations, the model with $P=2$ and $Q=0$ attained the highest average test accuracy of 99.95% $\pm$ 0.07%, outperforming all CNN-based counterparts. However, the best overall fault diagnosis performance was achieved by the PadéNet model with $P=2$, $Q=1$ and $\tanh$ activation, which reached an average test accuracy of 99.96% $\pm$ 0.04% and an F1-score of 99.96% $\pm$ 0.04%. A detailed summary of the test metrics for all 1D CNN, Self-ONN, and PadéNet models evaluated using the first accelerometer data is provided in Table 4.
Figure 5: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=2,Q=0$), and PadéNet ($P=2,Q=1$) models with the first accelerometer data as input. The 1D Self-ONN and PadéNet models use the $\tanh$ activation function, while the 1D CNN utilizes $\mathrm{LeakyReLU}$ with negative slope of 0.01.
@@
[…]
@@
Figure 6: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=2,Q=1$) models with the second accelerometer data as input. The 1D Self-ONN model uses the $\tanh$ activation function, while the 1D CNN and PadéNet utilize $\mathrm{LeakyReLU}$ with negative slope of 0.01.
Located on the bearing housing near the drive end, the second accelerometer presented a complex diagnostic environment due to potential signal disturbances from bearing dynamics. When the second accelerometer data was used as input, the 1D CNN model ($P=1$, $Q=0$) with $\tanh$ activation reached an average test accuracy of 93.94% $\pm$ 0.64% as given in Table 5. Switching to $\mathrm{LeakyReLU}$ activation across its convolutional layers significantly enhanced performance, boosting the average accuracy to 96.81% $\pm$ 0.51%, with reduced variability indicating improved consistency across each run. Among Self-ONN models tested, the configuration ($P=3$, $Q=0$) delivered the highest mean accuracy of 96.67% $\pm$ 0.34%. Although the average classification accuracy of the Self-ONN model ($P=3$, $Q=0$) surpassed that of the 1D CNN with $\tanh$ activation, switching to $\mathrm{LeakyReLU}$ enabled the 1D CNN to achieve higher fault diagnosis accuracy. Thus, it can be concluded that Self-ONNs generally outperform CNNs when $\tanh$ is employed as the activation function in both models. However, when $\mathrm{LeakyReLU}$ is used in all convolutional layers of a 1D CNN, it can achieve higher fault diagnosis accuracy compared to Self-ONNs utilizing the $\tanh$ activation function, potentially due to the vanishing gradient issue commonly associated with $\tanh$. In contrast to other 1D CNN and Self-ONN models, the most effective fault diagnosis for the second accelerometer was achieved by the 1D PadéNet model with $P=2,Q=1$ and $\mathrm{LeakyReLU}$ activation, having a mean test accuracy of 98.26% $\pm$ 0.26%, along with an F1-score of 98.26% $\pm$ 0.26%. A complete summary of performance metrics for all tested 1D CNN, Self-ONN, and PadéNet models using the second accelerometer data is available in Table 5.
@@
Figure 6 shows the aggregated confusion matrices across 5 independent runs for the 1D CNN $(P=1,Q=0)$, Self-ONN $(P=3,Q=0)$, and PadéNet $(P=2,Q=1)$ models, using data from the second accelerometer as input. For these confusion matrices, the 1D Self-ONN model uses the $\tanh$ activation function, while the 1D CNN and PadéNet utilize $\mathrm{LeakyReLU}$ with negative slope of 0.01. The 1D CNN $(P=1,Q=0)$ and Self-ONN $(P=3,Q=0)$ models achieve classification accuracies of 94.82% and 93.75%, respectively, in the faulty bearings (FB) class. In comparison, the 1D PadéNet $(P=2,Q=1)$ model delivers a significantly higher accuracy of 96.78% in the same class. Similarly, for the rotor misalignment (RM) class, the 1D CNN and Self-ONN models yield accuracies of 93.63% and 94.70%, respectively, whereas the 1D PadéNet $(P=2,Q=1)$ model once again outperforms them with a superior accuracy of 96.67%. Faulty bearings and rotor misalignment are two of the most critical fault types in electric motors, as they can severely compromise mechanical integrity and lead to costly operational downtimes if not detected early. Accurate classification of these faults is therefore essential for timely maintenance and fault prevention. The enhanced fault diagnosis performance of the 1D PadéNet model is clearly evident across all fault classes in these confusion matrices and reflects its superior capability to distinguish even the most critical fault types.
@@
[…]
Figure 7: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=1,Q=2$) models with the third accelerometer data as input. The 1D Self-ONN model uses the $\tanh$ activation function, while the 1D CNN and PadéNet utilize $\mathrm{LeakyReLU}$ with negative slope of 0.01.
When data from the third accelerometer, positioned on the bearing housing farthest from the drive end, is used as input, the results in Table 6 show that the 1D PadéNet model, particularly with $P=1$, $Q=2$ and $\mathrm{LeakyReLU}$ activation, consistently outperforms both the 1D CNN and Self-ONN models across all evaluated metrics. In particular, the 1D PadéNet model achieves the highest average test accuracy of 97.61% $\pm$ 0.27%, surpassing the best-performing configurations of the 1D CNN and Self-ONN models.
@@
[…]
Figure 8: Aggregated confusion matrices over 5 runs for the 1D CNN ($P=1,Q=0$), Self-ONN ($P=3,Q=0$), and PadéNet ($P=1,Q=2$) models with the acoustic data as input. The 1D Self-ONN model uses the $\tanh$ activation function, while the 1D CNN and PadéNet utilize $\mathrm{LeakyReLU}$ with negative slope of 0.01.
When acoustic signals captured by the microphone are used as input, 1D PadéNet models continue to demonstrate superior diagnostic capabilities, as shown in Table 7. The best performance was achieved by the configuration with $P=1,Q=2$ and $\mathrm{LeakyReLU}$ activation, which yielded the highest average accuracy of 98.33% $\pm$ 0.44%. This result is superior to those obtained using data from the second and third accelerometers, yet slightly underperforms compared to the first accelerometer, which is mounted directly above the drive-end.
@@
The fault diagnosis performance of 1D PadéNets is also compared with other DL-based methods trained on the University of Ottawa’s constant-speed vibration and audio datasets. In [10], a 15-layer 1D CNN-LSTM model (comprising 6 Conv1D and 2 LSTM layers) was proposed for classifying electrical and mechanical faults. The model takes an input window size of 1000 samples, as used in this study, with the same train-validation-test split ratios and all accelerometer inputs. To evaluate the impact of the LSTM layers, two other variations of the model were tested. First, the two LSTM layers were removed, resulting in a 13-layer CNN, referred to as “13-Layers CNN” in Table 8. In the second experiment, the LSTM layers in the 11th and 12th layers were replaced with Conv1D layers, producing a 15-layer CNN. All results are presented in Table 8 for each accelerometer sensor. For Accelerometer-1, the 1D PadéNet ($P=2$, $Q=1$) achieves the same average test accuracy as the 1D CNN-LSTM model but with significantly fewer trainable parameters, and outperforms all other CNN models. Moreover, for accelerometers 2 and 3, 1D PadéNet delivers comparable classification accuracy to the CNN-LSTM model, while the 13-layer and 15-layer CNN models fail to match those accuracies. However, it is worth noting that, due to the sequential nature of LSTM layers, the CNN-LSTM model tends to be slower during training and inference, whereas the 1D PadéNet achieves similar performance with greater computational efficiency.